OpenStudy (anonymous):

acceleration = 3t^2 + 2t + 1 find expression for instantaneous velocity at any time "t"

OpenStudy (anonymous):

I got the answer but here is integration constant included

OpenStudy (anonymous):

possibly. maybe there's more to this question than you're telling us :)

OpenStudy (anonymous):

no nothing more, i just wanted to know if integration constatnt should be there or not

OpenStudy (anonymous):

that's a stumper. Maybe just set it to zero?

OpenStudy (anonymous):

In the options it is given with and without a constant and the correct answer is the option without a constant

OpenStudy (anonymous):

set that to zero for what

OpenStudy (anonymous):

i don't want to find max or min

OpenStudy (anonymous):

i want to know then why integration constant is not included

OpenStudy (anonymous):

I meant set the constant to zero... thereby having pretty much no constant lawl

OpenStudy (anonymous):

why constant is 0

OpenStudy (anonymous):

I don't know... I'm just as confused as you are :D

OpenStudy (anonymous):

In Newtonian mechanics we don't assume constant

OpenStudy (anonymous):

Like in f=dp/dt f=kma when mass is constant k is 1 Newton so f= ma

OpenStudy (anonymous):

I don't know the reason

ganeshie8 (ganeshie8):

\[\large v(t) - v(0) = \int \limits_0^t 3t^2 + 2t + 1 ~dt\]

OpenStudy (anonymous):